# Efficient Backdoor Attacks for Deep Neural Networks in Real-world Scenarios

### v1

@@ 3.5 Entanglement Between Benign and Poisoning Features
In our data-constrained backdoor attacks, we have made two significant observations. Firstly, we have noticed a severe performance decline in class-constrained and domain-constrained backdoor attacks. Secondly, we consistently found that BadNets outperforms Blended in terms of attack efficiency. We attribute these observations to the entanglement between benign and poisoning features. Our study is the first to investigate feature entanglement in the context of backdoor attacks, providing new insights into backdoor learning. Ideally, we would expect backdoor models to rely solely on poisoning features to make decisions when encountering poisoning samples, as this would be the most efficient approach for backdoor attacks. However, neural networks tend to be greedy and utilize all features for decision-making [31], leading to activation of both poisoning and benign features during backdoor injection. This results in reduced poisoning efficiency when there is a difference in benign features between the backdoor injection and activation phases, as observed in data-constrained backdoor attacks.
We have further investigated our hypothesis and present our findings in Fig. 6. As shown in Fig. 6 (a), the attack efficiency of number-constrained, dirty-label single-class, and clean-label single-class backdoor attacks5.1 decreases in turn under the same poison rate. To understand the reason behind this phenomenon, we provide visualizations of the backdoor injection and activation phases for these three attacks in Fig. 6 (b). For the number-constrained backdoor attack, the distribution of poisoning samples (consisting of both benign and poisoning features) in the backdoor injection phase is the same as that in the backdoor activation phase. In other words, both benign and poisoning features are activated simultaneously during both phases. However, for the dirty-label single-class backdoor attack, the distribution of poisoning samples (consisting of single-class benign and poisoning features) in the backdoor injection phase is different from that in the backdoor activation phase. During the injection phase, both benign and poisoning features are activated, but during activation phase, only the poisoning feature is activated. This is the reason why previous attack methods on dirty-label single-class backdoor attacks exhibit performance degeneration. The clean-label single-class backdoor attack is similar to the dirty-label single-class backdoor attack in terms of the distribution of poisoning samples. However, during backdoor injection, there is competing activationIn the clean-label single-class backdoor attack, the benign feature of the accessible class (the same as the attack-target class) in both poisoning and clean sets is labeled with the same label (e.g., ”Fish” in Fig. 6), and the clean set contains more samples of the attack-target class. As a result, the presence of the benign feature in the poisoning set hampers the activation of the poisoning features. In contrary, the benign feature of the accessible class in poisoning and clean sets is labeled with the different label in the dirty-label single-class backdoor attack (e.g., the benign feature in the clean set is labeled as ”Frog”, while the benign+poisoning feature in the poisoning set is labeled as ”Fish”). Consequently, the benign feature in the poisoning set does not impact the activation of the poisoning features. between benign and poisoning features. Consequently, the poisoning efficiency of clean-label single-class backdoor attacks is lower than that of dirty-label single-class backdoor attacks.
In summary, performance degeneration in data-constrained backdoor attacks can be attributed to entanglement between benign and poisoning features. BadNets exhibits simpler triggers compared to Blended, leading to less entanglement and a higher attack success rate.
@@ 4 CLIP-guided Backdoor Attacks Method
In this section, we present our approach, which consists of two components: Clean Feature Suppression and Poisoning Feature Augmentation. These components are independent of each other and can be seamlessly combined. Specifically, Clean Feature Suppression can be implemented through CLIP-based Clean Feature Erasing (CLIP-CFE), and Poisoning Feature Augmentation has included two technologies: CLIP-based Universal Adversarial Perturbations (CLIP-UAP) and CLIP-based Contrastive Feature Augmentation (CLIP-CFA). By seamlessly integrating these two orthogonal aspects, our method presents a comprehensive and versatile solution that addresses all three types of data-constrained backdoor attacks. Before proposing the attack method, we first introduce threat model considered in our work.
4.1 Threat Model
Attack goal. The objective of our paper is aligned with popular backdoor attacks, as seen in previous studies [16, 24]. The attackers aim to activate a hidden trigger within the model by providing specific inputs, leading the model to produce incorrect results. Our attack strategy emphasizes three key properties: (i) Minimal side effects: The backdoor attacks should not adversely impact the accuracy of the model on benign inputs. (ii) Effective backdoor: The attack should have a high success rate across various datasets and models, ensuring its efficiency. (iii) Stealthy attack: The backdoor attacks should be inconspicuous and difficult to detect, maintaining their stealthiness. Our research aims to develop invigorative backdoor attacks that strike a balance between effectiveness and preserving the integrity of the model’s performance on legitimate inputs.
Attackers’ capabilities. Building upon previous studies [16], we make the assumption that the attackers possess the capability to control the training data. However, we further impose a stricter assumption in this work, stating that the attackers have control over only a portion of the training data. Consequently, we divide the attack scenario into three distinct tasks, each representing different capabilities of the attacker. These tasks include: (i) Number-constrained backdoor attacks, where the attacker has access to only a subset of the training data; (ii) Class-constrained backdoor attacks, where the attacker has access to only a subset of the classes in the training data; and (iii) Domain-constrained backdoor attacks, where the attacker has access to only a subset of the domains within the training data. By considering these various constraints, we provide a comprehensive analysis of backdoor attacks in different data-constrained scenarios.
@@ 4.2 Clean Feature Suppression
As described in Sec. 3, the effectiveness of data-constrained backdoor attacks is hindered due to the entanglement of benign and poisoning features during the backdoor injection phase. To address this challenge, we propose a solution called "clean feature suppression" in this section. The primary objective of this approach is to minimize the impact of benign features on the decision-making process, thus amplifying the significance of poisoning features.
@@ 4.2.1 CLIP-based Clean Feature Erasing
To achieve clean feature suppression, we can employ a feature extractor pre-trained on the entire training set (As shown in Clean Feature Erasing Noise). However, since our data-constrained backdoor attacks lack access to the complete training set, an alternative solution is required. Recent studies have shown that pre-trained CLIP [46] generates consistent and robust semantic representations across a wide range of (image, text) pairs, enabling impressive zero-shot classification performance comparable to supervised learning accuracy on challenging datasets like ImageNet (As shown in CLIP for Zero-shot Classification). Hence, we can utilize the pre-trained general model CLIP, which replaces§§§CLIP [46] is a general pre-trained model that has been released by OpenAI. It is pre-trained on 400 million image-text pairs harvested from the Web and can express a much wider set of visual concepts. the feature extractor trained on the entire training set, allowing us to achieve clean feature suppression (As shown in CLIP for Clean Feature Erasing).
Clean feature erasing noise. The technique of clean feature suppression aims to eliminate the clean information present in images by introducing optimized noise, denoted as $\delta$, which helps modify the input image to resemble the unbiased class. In accordance with the data-constrained backdoor attack pipeline outlined in Sec. 3, we assume that the chosen clean training dataset for generating the poisoning set consists of $P$ clean examples, denoted as $\mathcal{P^{\prime}}\subset X\times Y$ (where $\mathcal{P^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,P\}$). Here, $x_{i}\in X$ represents the inputs, $y_{i}\in Y=\{1,2,\cdots,C\}$ represents the labels, and $C$ denotes the total number of classes. We refer to the modified version as $\mathcal{P}_{e}=\{(x_{e,i},y_{i})|i=1,\cdots,P\}$, where $x_{e,i}=x_{i}+\delta_{i}$ represents the erased version of the training example $x_{i}\in\mathcal{P^{\prime}}$. The term $\delta_{i}\in\Delta$ denotes the "invisible" noise applied to achieve the erasing effect. The noise $\delta_{i}$ is subject to the constraint $||\delta_{i}||_{p}\leq\epsilon$, where $||\cdot||_{p}$ represents the $L_{p}$ norm, and $\epsilon$ is set to a small value to ensure the stealthiness of the backdoor attacks. Our objective in erasing the clean features is to ensure that the pre-trained feature extractor does not extract any meaningful information from the given images $x$. This is achieved by introducing customized and imperceptible noise, denoted as $\delta_{i}$. To be more specific, for a clean example $x_{i}$, we propose to generate the noise $\delta_{i}$ that erases the features by solving the following optimization problem:
@@
CLIP for zero-shot classification. The pre-trained CLIP model [46] possesses the ability to express a broader range of visual concepts and has been utilized as a general feature extractor in various tasks. These tasks include text-driven image manipulation [44], zero-shot classification [6], and domain generalization [42]. In this section, we introduce the pipeline of CLIP in zero-shot classification, which can serve as inspiration for incorporating it into our clean feature erasing approach. CLIP achieves zero-shot classification by aligning text and image features. Firstly, CLIP employs its text encoder, denoted as ^$\hat{\mathcal{E}}_{t}(\cdot)$, to embed the input prompts ("a photo of a $c_{i}$") into text features $T_{i}\in\mathbb{R}^{d}$, where $i=\{1,2,\cdots,C\}$ represents the classes. Subsequently, the image feature $I_{j}\in\mathbb{R}^{d}$ of image $x_{j}$ is embedded using the image encoder, denoted as ^$\hat{\mathcal{E}}_{i}(\cdot)$. During the inference phase, the classification prediction $y_{j}$ is computed using the cosine similarity between $T_{i}$ and $I_{j}$. This can be expressed as:
[…]
CLIP for Clean Feature Erasing (CLIP-CFE). Taking inspiration from CLIP’s approach to zero-shot classification, we leverage a general CLIP model to optimize the feature erasing noise. This allows us to relax the need for a proxy feature extractor pre-trained on the entire training set. We consider $C$ prompts, "a photo of a $c_{i}$," corresponding to different classes $c_{i}$ in the dataset, where $i=1,\cdots,C$. The CLIP-based feature erasing noise, denoted as $\delta_{i}$, is proposed for the input $x_{i}$ by solving the following optimization problem:
@@
$f_{CLIP}(x_{i}+\delta_{i},\mathbb{P})=$ (6)
[^$\bigg{[}\frac{\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{1})\rangle}{\sum_{i=1}^{C}\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{i})\rangle},\cdots,\frac{\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{C})\rangle}{\sum_{i=1}^{C}\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{i})\rangle}\bigg{]}.$
To solve the constrained minimization problem illustrated in Eq. 4, we utilize the first-order optimization method known as Projected Gradient Descent (PGD) [37]. The PGD method enables us to find a solution by iteratively updating the noise as follows:
@@ 4.3 Poisoning Feature Augmentation
In addition to eradicating clean features in images to tackle the entanglement between benign and poisoning features, enhancing the expression of poisoning features is another effective approach. In this section, we present two parallel triggers aimed at augmenting the poisoning features.
4.3.1 CLIP-based Universal Adversarial Perturbations
In this section, we also employ the widely-used pre-trained CLIP model [46] to generate universal adversarial perturbations as the backdoor trigger. Xia et al. [61] argue that deep models inherently possess flaws, and it is easier to exploit and enhance an existing flaw to serve as a backdoor rather than implanting a new one from scratch (BadNets [16] and Blended [5]). Universal Adversarial Perturbations (UAP) [70, 61] utilize these inherent flaws in models as triggers, providing a straightforward method for augmenting the poisoning feature. However, this approach typically requires a feature extractor that has been pre-trained on the entire training set, which is not practical in data-constrained backdoor attacks. To address this limitation, we propose a CLIP-based§4.2.1 Universal Adversarial Perturbations (CLIP-UAP) method. Specifically, given an accessible clean training set $\mathcal{D^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,N^{\prime}\}$ and an attack-target label $k$, the defined trigger can be formulated as follows:
where $\mathbb{P}$ and $f_{CLIP}$ are defined as shown in Eq. 5 and Eq. 6, respectively. Similar to Eq. 7, we utilize the first-order optimization method known as Projected Gradient Descent (PGD) [37] to solve the constrained minimization problem. The optimization process can be expressed as follows:
@@ 4.3.2 CLIP-based Contrastive Feature Augmentation
Similar to Eq. 10, we also adopt the first-order optimization method PGD [37] to solve the constrained minimization problem as follows:
@@ 4.4 Attack Summary
In Sec. 4.3, we present two independent trigger design methods: CLIP-based Universal Adversarial Perturbations (CLIP-UAP) and CLIP-based Contrastive Feature Augmentation (CLIP-CFA). These triggers are aimed at enhancing the expression of poisoning features and can replace previous trigger design approaches, leading to improved performance in data-constrained backdoor attacks. Additionally, in Sec. 4.2, we introduce a CLIP-based Clean Feature Erasing (CLIP-CFE) method. This approach minimizes the influence of clean features during the poisoning process and can be integrated into any of the aforementioned trigger design methods. By combining trigger design and clean feature erasing, our final approach achieves state-of-the-art performance in all three types of data-constrained backdoor attacks.
@@ 5 Experiments
We provides an overview of the experimental settings in this paper, covering various aspects such as datasets, model architecture, evaluation metrics, baselines, and implementations (Appendix 7.1). Subsequently, we perform comprehensive experiments to assess the effectiveness of our proposed methods through answering the following research questions:
RQ1: Are proposed technologies effective on three backdoor attacks? (Sec. 5.1)
RQ2: Are proposed technologies harmless for Benign Accuracy? (Sec. 5.2)
RQ3: Are proposed technologies stealthy for victims? (Sec. 5.3)
RQ4: Are proposed technologies effective for different poisoning settings? (Sec. 5.4)
In this section, we present the results specifically for CIFAR-100 datasets. Experimental outcomes for CIFAR-10 and ImageNet-50 are provided in Appendix 7.2 and Appendix 7.3 respectively. Additionally, for further discussions, please refer to Appendix 7.5.
@@ 5.1 RQ1: Are proposed technologies effective on three backdoor attacks?
To assess the effectiveness of our proposed technologies, we conduct attacks on various target models and datasets, evaluating the Attack Success Rate (ASR) for each target model. In order to establish a basis for comparison, we introduce two baseline attack methods: BadNets [16] and Blended [5], as discussed in Sec. 2.1.2. Fig. 7 illustrates the performance of the following types of backdoor attacks on the CIFAR-100 dataset: (a) number-constrained, (b) clean-label single-class (class-constrained), (c) dirty-label single-class (class-constrained), and (d) out-of-the-domain (domain-constrained)Both the clean-label single-class and dirty-label single-class backdoor attacks represent extreme scenarios of the class-constrained backdoor attack. In the clean-label single-class attack, the targeted class category is set to $Y^{\prime}={k}$, while in the dirty-label single-class attack, it is set to $Y^{\prime}={c}$ where $c\neq k$. Similarly, the out-of-the-domain backdoor attack is an extreme scenario of the domain-constrained backdoor attack, with a domain rate of 0. For further details, please refer to Appendix 7.1..
CLIP-based poisoning feature augmentation is more effective than previous attack methods. Our proposed methods, CLIP-UAP and CLIP-CFA, outperform the baseline techniques (BadNets [16] and Blended [5]While there are several more effective techniques for poisoning attacks, they typically necessitate access to the entire training data, rendering them unsuitable for our data-limited backdoor attacks.) in terms of consistency across different attacks and target models. Specifically, we achieved an ASR of 0.878, 0.825, 0.984, and 0.988 for BadNets, Blended, CLIP-UAP, and CLIP-CFA, respectively, in the number-constrained backdoor attack on the VGG-16 dataset. These results provide evidence that our proposed poisoning feature augmentation generates more effective triggers compared to other methods.
More discussion. While our technologies have shown significant improvements in poisoning efficiency compared to baseline methods, there are still important discussions that need to be addressed. We aim to provide answers to the following questions in a systematic manner in Appendix 7.5: i) Why do we observe performance degradation in the clean-label single-class attack? ii) Why are domain-constrained backdoor attacks generally easier compared to class-constrained backdoor attacks?

### v2

@@ 3.2 Entanglement Between Benign and Poisoning Features
In Sec A.3, we present three observations pertaining to data-constrained backdoor attacks, serving as compelling evidence for the presence of a significant interdependence between benign and poisoning features during the backdoor injection. This intricate entanglement is identified as the primary factor responsible for the inadequacies exhibited by current attack methodologies within data-constrained scenarios. Our study is a pioneering exploration of feature entanglement in the context of backdoor attacks, yielding fresh and valuable insights into the realm of backdoor learning. Ideally, one would anticipate backdoor models to exclusively rely on poisoning features when confronted with poisoning samples, as this would be the most efficient strategy for executing successful backdoor attacks. However, neural networks tend to be greedy and utilize all features for decision-making Li et al. (2023c), leading to activation of both poisoning and benign features during backdoor injection. This results in reduced poisoning efficiency when there is a difference in benign features between the backdoor injection and activation phases, as evidenced in data-constrained backdoor attacks.
@@ 4 CLIP-guided Backdoor Attacks Method
We present our approach, which consists of two components: Clean Feature Suppression and Poisoning Feature Augmentation. These components are independent of each other and can be seamlessly combined. The threat model considered in our study has been introduced in Sec. A.4
@@ 4.1 Clean Feature Suppression
As described in Sec. 3, the effectiveness of data-constrained backdoor attacks is hindered due to the entanglement of benign and poisoning features during the backdoor injection phase. To address this challenge, we propose a solution called ”clean feature suppression” in this section. The primary objective of this approach is to minimize the impact of benign features on the decision-making process, thus amplifying the significance of poisoning features.
@@ 4.1.1 CLIP-based Clean Feature Erasing
To achieve clean feature suppression, we can employ a feature extractor pre-trained on the entire training set (As shown in Sec. Clean Feature Erasing Noise). However, since our data-constrained backdoor attacks lack access to the complete training set, an alternative solution is required. Recent studies have shown that pre-trained CLIP Radford et al. (2021) generates consistent and robust semantic representations across a wide range of (image, text) pairs, enabling impressive zero-shot classification performance comparable to supervised learning accuracy on challenging datasets like ImageNet (As shown in A.5). Hence, we can utilize the pre-trained general model CLIP, which replaces the feature extractor trained on the entire training set, allowing us to achieve clean feature suppression (As shown in CLIP for Clean Feature Erasing).
Clean Feature Erasing Noise. The technique of clean feature suppression aims to eliminate the clean information present in images by introducing optimized noise, denoted as $\delta$, which helps modify the input image to resemble the unbiased class. In accordance with the data-constrained backdoor attack pipeline outlined in Sec. 3, we assume that the chosen clean training dataset for generating the poisoning set consists of $P$ clean examples, denoted as $\mathcal{P^{\prime}}\subset X\times Y$ (where $\mathcal{P^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,P\}$). Here, $x_{i}\in X$ represents the inputs, $y_{i}\in Y=\{1,2,\cdots,C\}$ represents the labels, and $C$ denotes the total number of classes. We refer to the modified version as $\mathcal{P}_{e}=\{(x_{e,i},y_{i})|i=1,\cdots,P\}$, where $x_{e,i}=x_{i}+\delta_{i}$ represents the erased version of the training example $x_{i}\in\mathcal{P^{\prime}}$. The term $\delta_{i}\in\Delta$ denotes the ”invisible” noise applied to achieve the erasing effect. The noise $\delta_{i}$ is subject to the constraint $||\delta_{i}||_{p}\leq\epsilon$, where $||\cdot||_{p}$ represents the $L_{p}$ norm, and $\epsilon$ is set to a small value to ensure the stealthiness of the backdoor attacks. Our objective in erasing the clean features is to ensure that the pre-trained feature extractor does not extract any meaningful information from the given images $x$. This is achieved by introducing customized and imperceptible noise, denoted as $\delta_{i}$. To be more specific, for a clean example $x_{i}$, we propose to generate the noise $\delta_{i}$ that erases the features by solving the following optimization problem:
@@
CLIP for Clean Feature Erasing (CLIP-CFE). Taking inspiration from CLIP’s approach to zero-shot classification (Sec. A.5), we leverage a general CLIP model to optimize the feature erasing noise. This allows us to relax the need for a proxy feature extractor pre-trained on the entire training set. We consider $C$ prompts, ”a photo of a $c_{i}$,” corresponding to different classes $c_{i}$ in the dataset, where $i=1,\cdots,C$. The CLIP-based feature erasing noise, denoted as $\delta_{i}$, is proposed for the input $x_{i}$ by solving the following optimization problem:
@@
$f_{CLIP}(x_{i}+\delta_{i},\mathbb{P})=\bigg{[}\frac{\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{1})\rangle}{\sum_{i=1}^{C}\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{i})\rangle},\cdots,\frac{\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{C})\rangle}{\sum_{i=1}^{C}\langle\hat{\mathcal{E}}_{i}(x_{i}+\delta_{i}),\hat{\mathcal{E}}_{t}(p_{i})\rangle}\bigg{]}.$ (5)
To solve the constrained minimization problem illustrated in Eq. 3, we utilize the first-order optimization method known as Projected Gradient Descent (PGD) Madry et al. (2017). The PGD method enables us to find a solution by iteratively updating the noise as follows:
@@ 4.2 Poisoning Feature Augmentation
In addition to eradicating clean features in images to tackle the entanglement between benign and poisoning features, enhancing the expression of poisoning features is another effective approach. In this section, we present two parallel triggers aimed at augmenting the poisoning features: CLIP-based Contrastive Feature Augmentation (4.2.1) and CLIP-based Universal Adversarial Perturbations (A.6).
@@ 4.2.1 CLIP-based Contrastive Feature Augmentation
Similar to Eq. 14, we also adopt the first-order optimization method PGD Madry et al. (2017) to solve the constrained minimization problem as follows:
@@ 4.3 Attack Summary
we present two independent trigger design methods: CLIP-UAP (Sec. A.6) and CLIP-CFA (Sec. 4.2.1). These triggers are aimed at enhancing the expression of poisoning features and can replace previous trigger design approaches, leading to improved performance in data-constrained backdoor attacks. Additionally, in Sec. 4.1, we introduce a CLIP-CFE method. This approach minimizes the influence of clean features during the poisoning process and can be integrated into any aforementioned trigger design methods. By combining trigger design and clean feature erasing, our final approach achieves state-of-the-art performance in all types of data-constrained backdoor attacks.
@@ 5 Experiments
We provides an overview of the experimental settings, covering datasets, model architecture, evaluation metrics, baselines, and implementations (Appendix A.7). Subsequently, we perform comprehensive experiments to assess the effectiveness of our proposed methods through answering the following research questions: RQ1: Are proposed technologies effective on three backdoor attacks? (Sec. 5.1). RQ2: Are proposed technologies harmless for Benign Accuracy? (Sec. 5.2). RQ3: Are proposed technologies stealthy for victims? (Sec. A.8). RQ4: Are proposed technologies effective for different poisoning settings? (Sec. 5.3). In this section, we present the results specifically for CIFAR-100 datasets. Experimental outcomes for CIFAR-10 and ImageNet-50 are provided in Appendix A.9 and Appendix A.10 respectively. Additionally, for more experiments and further discussions, please refer to Appendix A.11, A.12, A.13, A.14, and A.15.
To assess the effectiveness of our proposed technologies, we conduct attacks on various target models and datasets, evaluating the Attack Success Rate (ASR) for each target model. In order to establish a basis for comparison, we introduce two baseline attack methods: BadNets Gu et al. (2019) and Blended Chen et al. (2017), as discussed in Sec. A.2.1. Fig. 3 illustrates the performance of the following types of backdoor attacks on the CIFAR-100 dataset: (a) number-constrained, (b) clean-label single-class (class-constrained), (c) dirty-label single-class (class-constrained), and (d) out-of-the-domain (domain-constrained) [^1] Both the clean-label single-class and dirty-label single-class backdoor attacks represent extreme scenarios of the class-constrained backdoor attack. In the clean-label single-class attack, the targeted class category is set to $Y^{\prime}={k}$, while in the dirty-label single-class attack, it is set to $Y^{\prime}={c}$ where $c\neq k$. Similarly, the out-of-the-domain backdoor attack is an extreme scenario of the domain-constrained backdoor attack, with a domain rate of 0. For further details, please refer to Appendix A.7..
CLIP-based poisoning feature augmentation is more effective than previous attack methods. Our proposed methods, CLIP-UAP and CLIP-CFA, outperform the baseline techniques (BadNets Gu et al. (2019) and Blended Chen et al. (2017)  [^2] While there are several more effective techniques for poisoning attacks, they typically necessitate access to the entire training data, rendering them unsuitable for our data-limited backdoor attacks.) in terms of consistency across different attacks and target models. Specifically, we achieved an ASR of 0.878, 0.825, 0.984, and 0.988 for BadNets, Blended, CLIP-UAP, and CLIP-CFA, respectively, in the number-constrained backdoor attack on the VGG-16 model. These results provide evidence that our proposed poisoning feature augmentation generates more effective triggers compared to other methods.
More discussion. While our technologies have shown significant improvements in poisoning efficiency compared to baselines, there are still important discussions that need to be addressed. We aim to provide answers to the following questions in a systematic manner in Appendix A.15: i) Why do we observe performance degradation in the clean-label single-class attack? ii) Why are domain-constrained backdoor attacks generally easier compared to class-constrained backdoor attacks?